1.log11(x2+3x) =log11(3x+16)

2.In(4-3x) -In(7x) =In(11)

3.log(x) +log(x+12)=log(x-10)

4.In(x) =In(15-x) -In(x+1)

5.log8(4x+1) =-1

6.log6(3x) -log6(x+5) =1


1
Expert's answer
2021-03-24T14:03:37-0400

1.

"x^2+3x=3x+16, 3x+16>0"

"x^2=16, x>-\dfrac{16}{3}"

"x_1=-4, x_2=4"

2.

"\ln(4-3x)=\ln(77x), 0<x<\dfrac{4}{3}"

"4-3x=77x"

"x=0.05"


3.


"\log(x(x+12))=\log(x-10), x>10"

"x^2+12x=x-10"

"x^2+11x+10=0"

"(x+1)(x+10)=0"

Since "x<0," then there are no solutions.

There are no solutions.


4.


"\ln(x)+\ln(x+1)=\ln(15-x), 0<x<15"

"\ln(x(x+1))=\ln(15-x)"

"x^2+x=15-x"

"x^2+2x-15=0"




"(x+5)(x-3)=0"

Since "0<x<15," we take "x=3."

Solution is "x=3."


5.


"4x+1=8^{-1}"

"4x=-\dfrac{7}{8}"

"x=-\dfrac{7}{32}"

6.


"\log_6(3x)=\log_6(6(x+5)), x>0"

"3x=6(x+5)"

"x=2x+10"

Since "x=-10<0," then there are no solutions.

There are no solutions.